Samples of 3 cards are drawn from a population of five cards numbered from 1-5.


1. How many are the possible outcomes?


2. What are the possible means?


3. What is the probability of getting 4 as a mean?


4. What is the probability of getting 2 as a mean?


5. What is the probability of getting 3.33 as a mean?

1
Expert's answer
2022-03-19T02:39:57-0400

1.    There are "C_{5}^{3}=10" possible outcomes.

2.    The outcomes are

"\left( 1,2,3 \right) ,\bar{x}=2\\\left( 1,2,4 \right) ,\bar{x}=\frac{7}{3}\\\left( 1,2,5 \right) ,\bar{x}=\frac{8}{3}\\\left( 1,3,4 \right) ,\bar{x}=\frac{8}{3}\\\left( 1,3,5 \right) ,\bar{x}=3\\\left( 1,4,5 \right) ,\bar{x}=\frac{10}{3}\\\left( 2,3,4 \right) ,\bar{x}=3\\\left( 2,3,5 \right) ,\bar{x}=\frac{10}{3}\\\left( 2,4,5 \right) ,\bar{x}=\frac{11}{3}\\\left( 3,4,5 \right) ,\bar{x}=4"

The possible means are "2,\frac{7}{3},\frac{8}{3},3,\frac{10}{3},\frac{11}{3},4"

"3:\\P\left( \bar{x}=4 \right) =P\left( \left( 3,4,5 \right) \right) =\frac{1}{10}\\4:\\P\left( \bar{x}=2 \right) =P\left( \left( 1,2,3 \right) \right) =\frac{1}{10}\\5:\\P\left( \bar{x}=3.33 \right) =P\left( \bar{x}=\frac{10}{3} \right) =P\left( \left( 1,4,5 \right) \right) +P\left( \left( 2,3,5 \right) \right) =\frac{2}{10}=\frac{1}{5}\\"